MONT 100N – Modeling the Environment Information on Midterm Exam October 17, 2017

General Information and Groundrules

As announced in class and on the revised course schedule online, the midterm exam for our Montserrat seminar will be given in class on Friday, October 27.

• This will be a full period, individual exam. No sharing of information with others in any form will be permitted during the exam.

• You may use a calculator during the exam, but no other electronic devices.

• There will be three or four mathematical problems (each possibly with a few separate parts). These questions will be similar to things you have seen on the problem sets or the group projects from Chapters 1 - 4 of our text. Some sample exam questions are given later in this document.

• The exam will also include an essay question on a set topic chosen from the topics listed below. This part of the exam should take you about 15 minutes, so it will be necessary to spend some of your preparation time on deciding what you want to say for each topic.

Possible Essay Topics – I will choose one for the exam

1. In §193 of the Laudato Si’ encyclical letter, Pope Francis seems to suggest that economic growth should be be contained or reversed in order to achieve “sustainable development.” But economic growth in the developing world has lifted millions into the “global middle class” in recent decades and there is strong sentiment in countries like India that they deserve to experience the benefits of economic development like the development that has occurred in North America and Europe. Do more developed countries have the right to say “no” to those desires? Francis suggests redefining what we mean by “progress” to address these questions (§194), but is that realistic or desirable?

2. At the end of Chapter 1 of Laudato Si’, Francis says: “As often occurs in periods of deep crisis which require bold decisions, we are tempted to think that what is happening is not entirely clear. Superficially, apart from a few obvious signs of pollution and deterioration, things do not look that serious, and the planet could continue as it is for some time. Such evasiveness serves as a licence to carrying on with our present lifestyles and models of production and consumption. This is the way human beings contrive to feed their self-destructive vices: trying not to see them, trying not to acknowledge them, delaying the important decisions and pretending that nothing will happen.” Is this a fair characterization of the way things stand in the U.S. at present? What is Francis referring to with the “we are tempted to think that what is happening is not entirely clear?” Do powerful economic interests have a role in exaggerating and exploiting that uncertainty?

Sample Mathematical Questions

Disclaimer: The actual exam questions will cover the same mathematical topics and will be at approximately the same level of difficulty. However the organization and the content may be completely different.

A. (Refer to the tables in the text; something like those would be supplied for your use on a question like this.)

1. A committee on rail freight transport recommends increasing the mass that can be carried in a trailer from 42 to 46 tonnes (metric tons). How much are these masses in kilograms? In pounds?

2. About 24 million liters of rainwater are discharged by a storm drain into a river. The concentration of toxic arsenic in the rainwater averages 170 micrograms of arsenic per liter. How many kilograms of arsenic are discharged into the river each year.

B. In 2002, 10731 species were listed on the World Conservation Union “red list” of endangered species. In 2004, the corresponding number was 15042.

1. What was the ratio of the size of the 2004 list to the size of the 2002 list?

2. What was the percentage increase in the size of the list from 2002 to 2004?

3. It is not known exactly how many species of plants and animals there are on the Earth today. But say that number is estimated as 12 million. What percentage of that total number appeared on the red list in each year?

C. The following table gives several African countries, their total forest areas, and their areas of legally protected forests (in square kilometers)

Country  Total Protected
Cameroon 289965  17845
Congo  278979  12935
Gabon  239369  8975
Zaire  1439178 93160

1. Which country has the largest protected forest area?

2. Which country has the largest percentage of protected forest?

D. A cubic meter of air has a mass of about 1.25 kg. About 2 mg of that mass is methane. What is the mass concentration of methane in air, expressed in parts per million?

E. The US EPA Toxic Release Inventory for 2000 includes the following information:

State Toxics (in 10^{6}lb)
Nevada  1000
Utah  956
Arizona  744
Alaska  535
Texas  302

1. Construct a bar chart showing this information. Label your vertical axis.

2. Construct pie chart showing the portions of the total toxics from these five states that each state contributes.

3. Which state or states is approximately “twice as toxic” as Alaska.

F. Look at the plot in the spreadsheet link from the course homepage next to the link where you downloaded this review sheet. The data shows the gray wolf population in Wisconsin from 1980 to 2002.

1. Note that the portion of the graph representing the years 1980 to 1992 looks quite linear. Estimate a linear model for this period by using the line through the two points (1980, 50) and (1990, 55). (Determine the equation of the line through those points.)

2. Do the same for the period 1993-2002 using the two points (1993, 60) and (2002, 325).

3. What are the slopes of these two lines? What are the units of the slope and what do they represent in real world terms?

4. Look at the least squares regression line for the whole data set, together with the computed slope and intercept values for the regression line. Which method seems superior in this case – using two different lines for different time periods, or one regression line for the whole time? Explain how you are making this decision.